# Modeling Time-Variant Responses of Optical Compressors with Selective State Space Models

### v2

@@ 3.1 Network Architecture
The architecture we propose for modeling analog optical compression is centered on Selective State Space (S6) layers, as illustrated in Figure 1. It includes a linear FC layer that processes and compresses the input, which then feeds into an S6 block. The conditioning layers process this block’s output based on the values of the compression parameters $\boldsymbol{p}$, before passing its output through a second, identical S6 block. The output layer consists of a single unit that predicts a coefficient. When multiplied by the input sample, this coefficient produces the output sample. The architecture generates one output audio sample per inference cycle, taking as input a buffer of the $64$ most recent input samples. This implies that with each inference cycle, a new input sample enters the buffer, while the oldest is discarded. Therefore, the set of the $64$ most recent input samples constitutes the network input, which, after initial compression by the FC layer, is used to update the states in the S6 layers. This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time. Predicting one sample at a time helps minimize the audible artifacts commonly associated with machine-learning models of audio effects, which arise from slight amplitude mismatches at the boundaries of consecutive output segments.
@@
$h_{n}$ $=Ah_{n-1}+Bx_{n}$
$y_{n}$ $=Ch_{n}+Dx_{n}$ (1)
where $x$ and $y$ are the input and output samples at time $n$. A(N × N), B(N × 1), C(1 × N), and D(1 × 1) are complex or real-valued matrices expressing linear mappings between $x_{n}$ and $h_{n}$ the states of the system. The layer captures the long-range dependencies using the state matrix A. This matrix encodes all the past input history in $h_{n}$, finding a map from $x_{n}$ to a higher N-dimensional space $h_{n}$ that represents the compression of the input’s past. In the S6 layer, B and C are dependent on the input and computed using a linear FC layer, while A and D’s parameters are independent of the input and learned during the training process.
@@
The S6 layer is used inside a block, as in  (Gu & Dao,, 2023), illustrated in Figure 2. The first layer of the block is a linear FC with a number of units equal to double the length of the desired internal dimension, thereby expanding the block’s input size. The chosen internal dimension is $3$, corresponding to the output size of the S6 layer, selected to limit the total number of model parameters. The expanded vector is split into two equal-sized vectors, one passing the convolutional layer and, afterward, swish function  (Ramachandran et al.,, 2017) and finally the S6 layer; the other is element-wise multiplied by the output of the S6 layer, after the swish function as well. The swish function acts as a gating mechanism to regulate the flow of information and features a learnable parameter, providing smoother behavior than a ReLU. Finally, since S6 does not include nonlinear operators, the block output is processed through a nonlinear FC with $2$ units and the Gaussian Error Linear Unit (GELU)  (Hendrycks & Gimpel,, 2023) activation function. Figure  2 illustrates the described process, including the size of each layer, which can be described as follows:
@@
As detailed in Section 4.1, the LA-2A features a continuous parameter called ’peak reduction’, which defines the amount of compression, and a switch to toggle between limit and compression modes, which influences the amount of compression of the device. The attack and release times are not controllable and depend on the input signal. Therefore, for the LA-2A, $\boldsymbol{p_{co}}$ includes the ’peak reduction’ and the binary value of the mode switch, and $\boldsymbol{p_{ti}}$ includes only the vector $\boldsymbol{f}$. On the other hand, the CL 1B offers four controllable parameters: threshold, ratio, attack, and release time. However, the threshold and ratio influence the amount of compression, while the attack and release times affect the timing behavior of the device. Therefore, for the CL 1B, $\boldsymbol{p_{co}}$ includes the values of the threshold and ratio, while $\boldsymbol{p_{ti}}$ includes the values of the attack and release times.
@@
To evaluate the effectiveness of the proposed architecture in modeling analog optical compression, we designed three additional architectures that share the same conditioning process but feature different recurrent blocks, specifically based on LSTM, ED, and S4D, as illustrated in Figure 4. This approach allows us to compare their capabilities in modeling nonlinearities and encoding temporal information and to compare their performance when integrated with the same conditioning process. The S4D architecture uses the S4D layer, which is also described by Equation 3.1; however, in this case, the parameters B, C, and D are trainable and not functions of the input $\boldsymbol{x}$. As mentioned earlier, the principal difference with the S6 variant lies in this aspect. As the S6, the S4D layer incorporates a FC layer followed by a GELU activation function, as it does not inherently include nonlinearity. The ED architecture utilizes an encoder-decoder structure inspired by our previous work (Simionato & Fasciani,, 2023), which processes the input segment with a convolutional layer to establish the initial internal states for the first LSTM layer. Unlike the first, the second LSTM layer relies on its own internal states. The ED architecture also incorporates a linear FC layer following the LSTM layers. The S4D architecture employs a specialized SSM layer, which is an efficient variant that parameterizes its state matrix as a diagonal matrix  (Gu et al.,, 2022b). The LSTM architecture consists of an LSTM layer followed by a linear FC layer, embodying the most traditional approach. In addition to the architectures described above, we also use several previously proposed models as baselines for comparing performance. These models include LSTM-b (Steinmetz & Reiss,, 2022), ED-b (Simionato & Fasciani,, 2023), and TCN-b (Steinmetz & Reiss,, 2022). The ED-b model employs a CNN layer to process the input signal and a linear FC layer to process the control parameters in the encoder. The outputs of these two processes are then combined to initialize the states for the LSTM in the decoder. Following the LSTM is an FC layer with a sigmoid activation function. The LSTM-b architecture comprises an LSTM layer that receives input samples concatenated with the control parameters. Both the LSTM-b and ED-b baselines have an identical input, including $64$ samples. They are designed to have a number of trainable parameters comparable to that of the proposed model—with the LSTM-b having $13$ units and the ED-b model having $6$ units for each layer. In contrast, TCNs are subject to specific architectural constraints, such as the number of input and output samples that define the receptive field. For the TCN-b, we opted for the highest-performing model described in the literature without making any modifications. This model processes $72,00$ input samples to generate $57,600$ output samples at each inference iteration—quantities significantly larger than those associated with our proposed architectures.
@@ 3.2 Computational Cost and Latency
Our proposed architecture has been designed to minimize computational complexity during inference by limiting the total number of trainable parameters to approximately $1,000$. We applied the same parameter limit to the other architectures included in our comparison to ensure a fair evaluation across different designs. Slight deviations from this limit are permissible due to the use of significantly different layers in each case. Starting from the equations defining each layer’s operations, we estimated the computational cost of inference for our proposed architecture in terms of Floating Point Operations (FLOPs) per audio sample. Our calculations of FLOPs consider both multiplication and addition operations. For the activation functions Softsign, Swish, GELU, and hyperbolic tangent, we have counted $10$ FLOPs per sample, respectively.
@@ 4.1 Datasets
The LA-2A Leveling Amplifier is an analog limiter/compressor manufactured by Teletronix, whose optical gain reduction mechanism works similarly to the one in the CL 1B. The LA-2A differs fundamentally from the CL1B compressor: the attack and release times can not be manually adjusted by users with knobs, as is possible with the CL 1B. Instead, the LA-2A features a fixed average attack time of $10$ ms and a multi-stage release. The first stage lasts $0.06$ seconds, and the second is governed by the response time of the light-sensitive resistor, which varies depending on the intensity and duration of the light-emitting element’s illumination of the resistor. The duration of this second stage can range from $0.5$ to $5$ seconds. This implies that if the compression is heavy and/or the signal has been above the threshold for an extended period, the LA-2A’s release time will be prolonged. Consequently, the release times for the LA-2A cannot be known a priori, as they are highly dependent on the signal’s history. The device features an input gain control knob, a switch to toggle between limiter or compressor mode, and a peak reduction knob that adjusts the amount of compression. Both knobs have markings with values that range from 00 and $100$.
@@
For the software emulations, we created the dataset using a fully automated software tool that we specifically developed to generate datasets from audio effects (Fasciani et al.,, 2024b). This tool enables the easy generation of datasets with a much larger and more precisely controlled combination of parameters than is possible with hardware analog devices, where the parameters are not controllable via an electrical interface. The dataset from the Universal Audio LA-2A software emulation was generated considering $100$ equally spaced (EQS) values that cover the full range of the peak reduction in both limiter and compressor modes. Also, in this case, the input gain was set to $23$. For the Softube CL 1B emulation, which has four variable parameters, increasing the number of values selected for each parameter significantly raises the number of unique parameter combinations and, consequently, the size of the dataset. This increment may lead to severe computational challenges when training the models, potentially rendering the training process infeasible. Therefore, for the Softube CL 1B, we generated two separate datasets: the first, tagged as ’dynamic,’ includes a variable ratio and threshold with attack and release times knobs fixed at $25\%$, which should correspond to $70$ ms and $1$ second, respectively; the second, tagged as ’time’, includes variable attack and release times with the ratio knob fixed $50\%$, which should correspond to $6:1$, and the threshold at $-20$ dBu. In both Softube CL 1B datasets, the variable parameters consist of $20$ EQS values that cover the full range of each parameter. Also, in this case, the output gain was set to 00, the sidechain bus was turned off, and the time-constant circuit mode was turned to manual.
@@ 4.2 Training
All models included in the experiments are trained using the Mean Squared Error (MSE) as a loss function, which has been demonstrated to effectively capture the dynamic aspects of the output signal, as shown in (Simionato & Fasciani,, 2023). The trainings are conducted for a maximum of $200$ epochs, with an early stopping condition triggered if no reduction in validation loss is observed for $30$ consecutive epochs. Models are trained using the input-output recordings in the dataset at the original sampling rate of $48$ kHz. The batch size is set to $2400$ examples, each consisting of $64$ input audio samples, the values of the control parameters, and the single output audio sample. Considering the network architecture described in Section 3.1, in which a buffer of the $64$ most recent samples is used as input, training layers with states necessitates sequential input-output examples. Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $1$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features. This configuration considers the $64$ input samples as the vector of input features at the current sampling instant. This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time. The control parameter values are scaled to the range [00, $1$], corresponding to the full range of the device. We use the Adam (Kingma & Ba,, 2014) optimizer with a gradient norm scaling of $1$  (Pascanu et al.,, 2013), and we designed an exponential decay learning rate ($lr$) as follows:
$lr=0.25^{e}*LR,\text{ with }LR=3\cdot 10^{4}$ (7)
@@ 4.3 Experiments
The proposed model, which incorporates S6 layers into its architecture, is evaluated against various alternative and baseline architectures, including those featuring LSTM, ED, S4D, and TCN layers. These models are trained on a dataset obtained from the analog optical dynamic range compressors Teletronix LA-2A and TubeTech CL 1B. Trainings use $90\%$ of the dataset, while the remaining $10\%$ is reserved for testing. The $90-10\%$ split is implemented at the individual recording level. Each recording represents the response for a unique combination of control parameters, which remain fixed during the recording process. This splitting strategy ensures that all combinations of control parameters are equally present in both the training and test sets, with the aim of developing an accurate model that replicates the behavior of the actual analog device. The initial $90\%$ of each recording is used for the train set, and the remaining portion is allocated to the test set. Manual adjustments are occasionally made to ensure that the splitting points occur during brief intervals of silence between sound sequences in the dataset. Modeling accuracy is assessed using the test set loss and additional quantitative metrics outlined in Section 4.3.1, and through listening tests (excluding the baseline models), as explained in Section 4.3.2.
@@
In previously described experiments, the models were evaluated on their ability to compress audio signals unseen during training but for control parameter combinations that were encountered during training. We conducted an additional experiment to assess the model’s capability to accurately compress audio signals for control parameter combinations not seen during training, thereby testing the models’ capacity for generalization in this respect. We also evaluated the impact of dataset size, specifically regarding the number and density of sampled values within the range of each control parameter. For these evaluations, we utilized the three datasets collected from the software emulations of Universal Audio LA-2A and Softube CL 1B. Their larger sizes and higher parameter sampling densities make it feasible to experiment with various subsets extrapolated from the datasets. These experiments involve three separate datasets: Universal Audio LA-2A, Softube CL 1B ’dynamic’, and Softube CL 1B ’time’. Each dataset is used three times with a fixed test set, while the size of the training set is varied. This results in a total of nine experimental settings, each used to train the various models involved in the comparative study. The test set comprises recordings, each with a duration of $195$ seconds, associated with combinations of control parameters that have normalized values (within the unit range) of $0.25$, $0.55$, or $0.75$. For the Universal Audio LA-2A dataset, this results in a total of $6$ combinations, with $3$ levels of peak reduction for each mode switch position. For the Softube CL 1B ’dynamic’ and ’time’ datasets, this yields $9$ combinations for each dataset, reflecting the two variable control parameters. The remaining portions of these datasets are used to create three separate training sets: the first includes all remaining recordings; for the second and third, we progressively reduce the number of EQS values selected for each variable parameter. For the Universal Audio LA-2A dataset, the reduction is from $100$ down to $20$ and $10$. For the Softube CL 1B ’dynamic’ and ’time’ datasets, the reduction goes from $20$ down to $10$ and $5$. A summary of these training and test sets is provided in Table 3.
@@
To further assess how well the models capture the time-variant response of the compressors and learn to emulate the time-behavior variations caused by the devices’ control parameters, we conducted an additional set of experiments. These experiments exclusively utilized TubeTech CL 1B recordings because this device allows for explicit compression attack and release times adjustment. In contrast, the LA 2A’s release time is influenced by the input signal. We selected four compression cases from the TubeTech CL 1B dataset, with the threshold set to $-10$ dBu and the ratio fixed at $6:1$. These settings result in discernible compression that is clearly noticeable but not extreme. For each case, the attack and release times differ, spanning from the fastest attack time of $0.5$ ms to the slowest of $500$ ms and from the quickest release time of $0.005$ seconds to the longest of $5$ seconds. We excluded the dataset’s slowest release setting of $10$ seconds since it is less commonly used in practical applications and is less suitable for assessing accuracy with short sound samples, as we do in this experiment. In the four selected settings, CL 1B presents a very diverse time-variant response, as the gain reduction is applied and removed with significantly different temporal patterns. We train all models, including the S6, alternative, and baseline architectures, on an exclusive subset of the dataset corresponding to one of the four selected compression settings. This approach aims to determine how well each model can learn specific temporal compression patterns and identify whether these diverse cases pose greater challenges. For testing, we utilize a short, $3.60$-second audio snippet that is not used for training and features bass and drum sounds. This compressed snippet has pronounced differences with the four distinct settings, and any modeling inaccuracies in compression attack and/or release are evident through the loss function and quantitative evaluation metrics. Finally, we train each model using all four compression settings to gauge the models’ ability to handle multiple parametric time-variant responses. This allows us to observe any degradation in the metrics that occurs when the model is conditioned with the selected compression settings. Given the small size of the training dataset and quicker convergence compared to previous experiments, we repeated each training 20 times. We report and analyze the performance metrics based on the average of these training iterations.

### v3

@@ 3.1 Network Architecture
The architecture we propose for modeling analog optical compression is centered on Selective State Space (S6) layers, as illustrated in Figure 1. It includes a linear FC layer that processes and compresses the input, which then feeds into an S6 block. The conditioning layers process this block’s output based on the values of the compression parameters $\boldsymbol{p}$, before passing its output through a second, identical S6 block. The output layer consists of a single unit that predicts a coefficient. When multiplied by the input sample, this coefficient produces the output sample.
The architecture generates one output audio sample per inference cycle, taking as input a buffer of the $64$ most recent input samples. This implies that with each inference cycle, a new input sample enters the buffer, while the oldest is discarded. Therefore, the set of the $64$ most recent input samples constitutes the network input, which, after initial compression by the FC layer, is used to update the states in the S6 layers. This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time. Predicting one sample at a time helps minimize the audible artifacts commonly associated with machine-learning models of audio effects, which arise from slight amplitude mismatches at the boundaries of consecutive output segments.
@@
$\boldsymbol{h}_{n}$ $=\boldsymbol{A}\boldsymbol{h}_{n-1}+\boldsymbol{B}\boldsymbol{j}_{n}$
$\boldsymbol{o}_{n}$ $=\boldsymbol{C}\boldsymbol{h}_{n}+\boldsymbol{D}\boldsymbol{j}_{n}$ (1)
where $\boldsymbol{j}_{n}$ and $\boldsymbol{o}_{n}$ are the input and output vectors at time $n$. $\boldsymbol{A}_{N\times N}$, $\boldsymbol{B}_{N\times M}$, $\boldsymbol{C}_{M\times N}$, and $\boldsymbol{D}_{M\times M}$ are complex or real-valued matrices expressing linear mappings between $\boldsymbol{j}_{n}$ and $\boldsymbol{h}_{n}$ the states of the system. The layer captures the long-range dependencies using the state matrix $\boldsymbol{A}$. This matrix encodes all the past input history in $\boldsymbol{h}_{n}$, finding a map from an M-dimensional space $\boldsymbol{j}_{n}$ to a N-dimensional space $\boldsymbol{h}_{n}$ that represents the compression of the input’s past. In the S6 layer, $\boldsymbol{B}$ and $\boldsymbol{C}$ are dependent on the input and computed using a linear FC layer, while $\boldsymbol{A}$ and $\boldsymbol{D}$’s parameters are independent of the input and learned during the training process.
@@
The S6 layer is used inside a block, as in  (Gu & Dao,, 2023), illustrated in Figure 2. The first layer of the block is a linear FC with a number of units equal to double the length of the desired internal dimension, thereby expanding the block’s input size. The chosen internal dimension is $3$, corresponding to the output size of the S6 layer, selected to limit the total number of model parameters. The expanded vector is split into two equal-sized vectors, one passing the convolutional layer and, afterward, swish function  (Ramachandran et al.,, 2017) and finally the S6 layer; the other is element-wise multiplied by the output of the S6 layer, after the swish function as well. The swish function acts as a gating mechanism to regulate the flow of information and features a learnable parameter, providing smoother behavior than a ReLU.
Finally, since S6 does not include nonlinear operators, the block output is processed through a nonlinear FC with $2$ units and the Gaussian Error Linear Unit (GELU)  (Hendrycks & Gimpel,, 2023) activation function. Swish and GELU activation functions, with continuous derivatives, have been preferred because they have shown empirically better performance than ReLU  (Ramachandran et al.,, 2017; Hendrycks & Gimpel,, 2023), a choice also made in  (Gu et al.,, 2022b; Gu & Dao,, 2023). Figure  2 illustrates the described process, including the size of each layer, which can be described as follows:
@@
As detailed in Section 4.1, the LA-2A features a continuous parameter called ‘peak reduction’, which defines the amount of compression, and a switch to toggle between limit and compression modes, which influences the amount of compression of the device. The attack and release times are not controllable and depend on the input signal. Therefore, for the LA-2A, $\boldsymbol{p_{co}}$ includes the ‘peak reduction’ and the binary value of the mode switch, and $\boldsymbol{p_{ti}}$ includes only the vector $\boldsymbol{f}$. On the other hand, the CL 1B offers four controllable parameters: threshold, ratio, attack, and release time. However, the threshold and ratio influence the amount of compression, while the attack and release times affect the timing behavior of the device. Therefore, for the CL 1B, $\boldsymbol{p_{co}}$ includes the values of the threshold and ratio, while $\boldsymbol{p_{ti}}$ includes the values of the attack and release times.
@@
To evaluate the effectiveness of the proposed architecture in modeling analog optical compression, we designed three additional architectures that share the same conditioning process but feature different recurrent blocks, specifically based on LSTM, ED, and S4D, as illustrated in Figure 4. This approach allows us to compare their capabilities in modeling nonlinearities and encoding temporal information and to compare their performance when integrated with the same conditioning process.
The S4D architecture uses the S4D layer, which is also described by Equation 3.1; however, in this case, the $\boldsymbol{B}$, $\boldsymbol{C}$, and $\boldsymbol{D}$ are independent of the input and learned during the training process. As mentioned earlier, the principal difference with the S6 variant lies in this aspect. As the S6, the S4D layer incorporates a FC layer followed by a GELU activation function, as it does not inherently include nonlinearity. The ED architecture utilizes an encoder-decoder structure inspired by our previous work (Simionato & Fasciani,, 2023), which processes the input segment with a convolutional layer to establish the initial internal states for the first LSTM layer. Unlike the first, the second LSTM layer relies on its own internal states. The ED architecture also incorporates a linear FC layer following the LSTM layers. The S4D architecture employs a specialized SSM layer, which is an efficient variant that parameterizes its state matrix as a diagonal matrix  (Gu et al.,, 2022b). The LSTM architecture consists of an LSTM layer followed by a linear FC layer, embodying the most traditional approach.
In addition to the architectures described above, we also use several previously proposed models as baselines for comparing performance. These models include LSTM-b (Steinmetz & Reiss,, 2022), ED-b (Simionato & Fasciani,, 2023), and TCN-b (Steinmetz & Reiss,, 2022). Both LSTM-b and ED-b baselines have an identical input, including $64$ samples. They are designed to have a number of trainable parameters comparable to that of the proposed model—with the LSTM-b having $13$ units and the ED-b model having $6$ units for each layer. In contrast, TCNs are subject to specific architectural constraints, such as the number of input and output samples that define the receptive field. For the TCN-b, we opted for the highest-performing model described in the literature without making any modifications. This model processes $7,200$ input samples to generate $57,600$ output samples at each inference iteration—quantities significantly larger than those associated with our proposed architectures.
@@ 3.2 Computational Cost and Latency
Our proposed architecture has been designed to achieve minimal computational complexity during inference by limiting the total number of trainable parameters to approximately $1,000$. This constraint has enabled us to reduce the computational load required to process a stereo audio signal at $48$ kHz to just under $200$ MFLOPS (Floating Point Operations Per Second). Such a limit is well-suited for most applications, including running multiple software audio effects concurrently on consumer-grade personal computers. Although implementations on dedicated DSP hardware can generally handle higher loads, we hypothesize that expanding the model size within the same architectural framework could yield more accurate emulations. Furthermore, the deliberate choice of relatively small models has facilitated the conducting of a substantial number of experiments involving architectural, dataset, and training variations, as the number of trainable parameters significantly impacts the training time.
To ensure a fair comparison in our study, we applied the same limit on the number of trainable parameters to the other architectures evaluated. We permitted minor deviations from this limit to accommodate the use of distinct layer types in each case. We have calculated the computational cost of inference for our proposed architecture in terms of Floating Point Operations (FLOPs) per audio sample by starting from the equations defining each layer’s operations. Our FLOP calculations include both multiplication and addition operations. For the Softsign, Swish, GELU, and hyperbolic tangent activation functions, we have estimated for $10$ FLOPs.
@@ 4.1 Datasets
The LA-2A Leveling Amplifier, an analog limiter/compressor manufactured by Teletronix, shares the optical gain reduction feature with the CL 1B but has distinct operational characteristics. Unlike the CL 1B, the LA-2A’s attack and release times are not adjustable; it has a fixed average attack time of $10$ ms and a two-stage release. The first stage lasts $0.06$ seconds, followed by a variable second stage determined by the light-sensitive resistor’s response, which can extend from $0.5$ to $5$ seconds. This variability means the LA-2A’s release time is signal-dependent and can lengthen if the compression is heavy or the signal consistently exceeds the threshold. Consequently, the release time for the LA-2A cannot be known a priori, as it is highly dependent on the signal’s history. The device features an input gain control knob, a switch to toggle between limiter or compressor mode, and a peak reduction knob that adjusts the amount of compression. Both knobs have markings with values that range from 00 and $100$.
@@
For the software emulations, we created the dataset using a fully automated software tool that we specifically developed to generate datasets from audio effects  (Fasciani et al.,, 2024b). This tool enables the easy generation of datasets with a much larger and more precisely controlled combination of parameters than is possible with hardware analog devices, where the parameters are not controllable via an electrical interface. The dataset from the Universal Audio LA-2A software emulation was generated considering $100$ equally spaced (EQS) values that cover the full range of the peak reduction in both limiter and compressor modes. Also, in this case, the input gain was set to $23$.
For the Softube CL 1B emulation, which has four variable parameters, increasing the number of values selected for each parameter significantly raises the number of unique parameter combinations and, consequently, the size of the dataset. This increment may lead to severe computational challenges when training the models, potentially rendering the training process infeasible. Therefore, for the Softube CL 1B, we generated two separate datasets: the first, tagged as ‘dynamic’, includes a variable ratio and threshold with attack and release times knobs fixed at $25\%$, which should correspond to $70$ ms and $1$ second, respectively; the second, tagged as ‘time’, includes variable attack and release times with the ratio knob fixed $50\%$, which should correspond to $6:1$, and the threshold at $-20$ dBu. In both Softube CL 1B datasets, the variable parameters consist of $20$ EQS values that cover the full range of each parameter. Also, in this case, the output gain was set to 00, the sidechain bus was turned off, and the time-constant circuit mode was turned to manual.
@@ 4.2 Training
All models included in the experiments are trained using the Mean Squared Error (MSE) as a loss function, which has been demonstrated to effectively capture the dynamic aspects of the output signal, as shown in (Simionato & Fasciani,, 2023). The trainings are conducted for a maximum of $200$ epochs, with an early stopping condition triggered if no reduction in validation loss is observed for $30$ consecutive epochs. Models are trained using the input-output recordings in the dataset at the original sampling rate of $48$ kHz. The batch size is set to $2400$ examples, each consisting of $64$ input audio samples, the values of the control parameters, and the single output audio sample.
Considering the network architecture described in Section 3.1, in which a buffer of the $64$ most recent samples is used as input, training layers with states necessitates an internal sequential order among the input-output examples. Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $1$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features. This configuration considers the $64$ input samples as the vector of input features at the current sampling instant. This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time. In contrast, a configuration with a sequence length of $64$ leads to more computations for the state updates, exacerbating the vanishing gradient problem and increasing the computational cost and training time of the network.
The control parameter values are scaled to the range [00, $1$], corresponding to the full range of the device. We use the Adam (Kingma & Ba,, 2014) optimizer with a gradient norm scaling of $1$  (Pascanu et al.,, 2013), and we designed an exponential decay learning rate ($lr$) as follows:
$lr=0.25^{e}*LR,\text{ with }LR=3\cdot 10^{-4}$ (7)
@@ 4.3 Experiments
The proposed model, which incorporates S6 layers into its architecture, is evaluated against various alternative and baseline architectures, including those featuring LSTM, ED, S4D, and TCN layers. These models are trained on a dataset obtained from the analog optical dynamic range compressors Teletronix LA-2A and TubeTech CL 1B. Trainings use $90\%$ of the dataset, while the remaining $10\%$ is reserved for testing. The $90-10\%$ split is implemented at the individual recording level. Each recording represents the response for a unique combination of control parameters, which remain fixed during the recording process. This splitting strategy ensures that all combinations of control parameters are equally present in both the training and test sets, with the aim of developing an accurate model that replicates the behavior of the actual analog device.
The initial $90\%$ of each recording is used for the train set, and the remaining portion is allocated to the test set. Manual adjustments are occasionally made to ensure that the splitting points occur during brief intervals of silence between sound sequences in the dataset. Modeling accuracy is assessed using the test set loss and additional quantitative metrics outlined in Section 4.3.1, and through listening tests (excluding the baseline models), as explained in Section 4.3.2.
@@
In previously described experiments, models were evaluated on their ability to compress audio signals unseen during training but for control parameter combinations encountered during training. We conducted additional experiments to assess the model’s capability to compress signals for control parameter combinations not seen during training. We also evaluated the impact of dataset size, specifically regarding the density of sampled values within the range of each control parameter. For these evaluations, we utilized the three datasets collected from the software emulations Universal Audio LA-2A and Softube CL 1B. Their larger sizes and higher parameter sampling densities allow for experimentation with various subsets of the datasets. The testing framework utilized three datasets: Universal Audio LA-2A, Softube CL 1B ‘dynamic’, and Softube CL 1B ‘time’. Each dataset is used three times with a fixed test set, while the size of the training is varied. This results in a total of nine comparative experiments. The test set comprises recordings, each with a duration of $195$ seconds, associated with combinations of control parameters that have normalized values of $0.25$, $0.55$, or $0.75$ (within the unit range).
For the Universal Audio LA-2A dataset, this results in a total of $6$ combinations, with $3$ levels of peak reduction for each mode switch position. For the Softube CL 1B ‘dynamic’ and ‘time’ datasets, this yields $9$ combinations for each dataset, reflecting the two variable control parameters. The remaining portions of these datasets are used to create three separate training sets: the first includes all remaining recordings; for the second and third, we progressively reduce the number of EQS values selected for each variable parameter. For the Universal Audio LA-2A dataset, the reduction is from $100$ down to $20$ and $10$. For the Softube CL 1B ‘dynamic’ and ‘time’ datasets, the reduction goes from $20$ down to $10$ and $5$. A summary of these training and test sets is provided in Table 3.
@@
To further assess how well the models capture the time-variant response of the compressors and learn to emulate the time-behavior variations caused by the devices’ control parameters, we conducted an additional set of experiments. These experiments exclusively utilized TubeTech CL 1B recordings because this device allows for explicit compression attack and release times adjustment. In contrast, the LA 2A’s release time is influenced by the input signal. We selected four compression cases from the TubeTech CL 1B dataset, with the threshold set to $-10$ dBu and the ratio fixed at $6:1$. These settings result in discernible compression that is clearly noticeable but not extreme. For each case, the attack and release times differ, spanning from the fastest attack time of $0.5$ ms to the slowest of $500$ ms and from the quickest release time of $0.005$ seconds to the longest of $5$ seconds. We excluded the dataset’s slowest release setting of $10$ seconds since it is less commonly used in practical applications and is less suitable for assessing accuracy with short sound samples, as we do in this experiment.
In the four selected settings, CL 1B presents a very diverse time-variant response, as the gain reduction is applied and removed with significantly different temporal patterns. We train all models, including the S6, alternative, and baseline architectures, on an exclusive subset of the dataset corresponding to one of the four selected compression settings. This approach aims to determine how well each model can learn specific temporal compression patterns and identify whether these diverse cases pose greater challenges. For testing, we utilize a short, $3.60$-second audio snippet that is not used for training and features bass and drum sounds. This compressed snippet has pronounced differences with the four distinct settings, and any modeling inaccuracies in compression attack and/or release are evident through the loss function and quantitative evaluation metrics.
Finally, we train each model using all four compression settings to gauge the models’ ability to handle multiple parametric time-variant responses. This allows us to observe any degradation in the metrics that occurs when the model is conditioned with the selected compression settings. Given the small size of the training dataset and quicker convergence compared to previous experiments, we repeated each training 20 times. We report and analyze the performance metrics based on the average of these training iterations.
Experiments were conducted on workstations equipped with an Intel Core i9-14900KF CPU, 192GB of DDR5 RAM at 5200MHz, and a GeForce RTX 4090 GPU. The models were implemented in Python using Tensorflow  [^5] https://www.tensorflow.org/. Training times varied across datasets and models, which are summarized in Table 4.
@@
Average Training Time per Epoch (min)
Set S6 S4D ED LSTM ED-b LSTM-b TCN-b
Teletronix LA-2A $7$ $7$ $7$ $12$ $4$ $4$ $1$
TubeTech CL 1B $37$ $35$ $34$ $35$ $23$ $22$ $3$
Universal Audio LA-2A-0.01 $25$ $23$ $22$ $36$ $24$ $24$ $1$
Universal Audio LA-2A-0.05 $5$ $5$ $5$ $5$ $6$ $5$ $1$
Universal Audio LA-2A-0.10 $2$ $5$ $4$ $4$ $2$ $2$ $1$
Softube CL 1B dynamic-0.05 $51$ $51$ $59$ $54$ $48$ $45$ $5$
Softube CL 1B dynamic-0.10 $27$ $27$ $26$ $25$ $17$ $16$ $1$
Softube CL 1B dynamic-0.20 $7$ $7$ $15$ $15$ $11$ $10$ $1$
Softube CL 1B time-0.05 $51$ $51$ $59$ $54$ $48$ $45$ $5$
Softube CL 1B time-0.10 $26$ $27$ $25$ $25$ $17$ $18$ $1$
Softube CL 1B time-0.20 $7$ $7$ $14$ $14$ $12$ $12$ $1$
Table 4: Average training time per epoch, including validation, across different datasets. Times are expressed in minutes.